OpenStudy (anonymous):

I need help quick please! Which of the following are solutions to the equation sinxcosx=sqrt(3)/4? 1. pi/12 +npi 2. pi/6 +npi 3. pi/12 +npi/2 4. 2pi/6 +npi I think the answers are 2&3.

OpenStudy (freckles):

recall sin(x)cos(x)=.5*2sin(x)cos(x)=.5sin(2x)

OpenStudy (anonymous):

So am I right with 2&3

OpenStudy (freckles):

How did you get 3?

OpenStudy (anonymous):

i dont know can you help me?

OpenStudy (akashdeepdeb):

\[\sin x.\cos x = \frac{\sqrt{3}}{4}\] \[\frac{1}{2} 2 \sin x.\cos x = \frac{\sqrt{3}}{4}\] \[\sin 2x = \frac{\sqrt{3}}{2}\] \[x = \frac{1}{2} \sin^{-1} (\frac{\sqrt{3}}{2})\]